### v2

# Classical spin liquids from frustrated Ising models in hyperbolic space

Abstract
Antiferromagnetic Ising models on frustrated lattices can realize classical spin liquids, with highly degenerate ground states and, possibly, fractionalized excitations and emergent gauge fields. Motivated by the recent interest in many-body system in negatively curved space, we study hyperbolic frustrated Ising models. Specifically, we consider nearest-neighbor Ising models on tesselations with odd-length loops in two-dimensional hyperbolic space. For finite systems with open boundaries we determine the ground-state degeneracy exactly, and we perform extensive finite-temperature Monte-Carlo simulations to obtain thermodynamic data as well as correlation functions. We show that the shape of the boundary, constituting an extensive part of the system, can be used to control low-energy states: Depending on the boundary, we find ordered or disordered ground states. Our results demonstrate how geometric frustration acts in curved space to produce classical spin liquids.
The behavior of many-body systems is crucially influenced by symmetry and geometry of the underlying real-space lattice, with band degeneracies, band topology, as well as frustration and related interference effects being key examples. Interest in recent years has extended to lattices embedded in curved space, hyperbolic space in particular, for a number of reasons: First, many condensed-matter paradigms established for Euclidean flat space are substantially modified in hyperbolic space, for instance band theory [1, 2, 3] and critical behavior [4, 5, 6]. Second, the holographic principle as realized by the AdS/CFT correspondence postulates a connection between interacting quantum field theories and gravity theories in higher dimensions [8, 7, 9]. As a generalization, it has also been postulated that gravity theories discretized on hyperbolic tilings of two-dimensional hyperbolic space are dual to spin chains at the boundary [10, 11, 12, 13, 14], a program known as ‘discrete holography’. Third, hyperbolic systems have been discussed in quantum information theory as candidates to efficiently encode qubits using toric code [15].
This paper is concerned with lattice models of interacting local moments in hyperbolic space. Defining such models has multiple aspects which differ from flat space: (i) The lattice geometry is dictated by the possibilities to discretize curved space in a periodic manner: This leads to distinct hyperbolic tilings [16] leading to distinct lattice geometries. (ii) Boundaries and boundary conditions can influence bulk properties even in the large-system limit. (iii) The form of magnetic interactions can be influenced by the curvature of the underlying space; this aspect appears particularly relevant for vectorial (as opposed to Ising-type) local moments [17].
Previous work on hyperbolic magnetism has studied the two-dimensional ferromagnetic Ising model in some detail [18, 19, 20, 6, 21], which shows a mean-field-type transition into a low-temperature ferromagnetic phase. Other papers have studied the Ising model with competing second-neighbor interaction [22] as well as with random-bond disorder [23]; and curvature as source of uniform frustration in continuum models has been discussed [24, 5].
In this paper, we focus on a different variety of hyperbolic magnetism, namely non-disordered frustrated lattice models. Given that frustration tends to suppress magnetic order, frustrated magnets herald many non-trivial forms of order and disorder. Prominent examples are various kinds of spin liquids, both classical and quantum, which often are characterized by fractionalized excitations and emergent gauge theories [25, 26, 27]. Notably, recent work has proposed realizations of the exactly solvable Kitaev model – where spin liquidity is driven by exchange frustration – on hyperbolic lattices [29, 28]. However, spin liquids driven by geometric frustration have not been considered to our knowledge, with the exception of a dimer-based wavefunction study in Ref. 30. We therefore choose to investigate the antiferromagnetic nearest-neighbor (NN) Ising model
$\mathcal{H}=J\sum_{\langle ij\rangle}\sigma_{i}\sigma_{j}$ (1)
where $\sigma_{i}=\pm 1$ corresponds to spin-$1/2$ degrees of freedom.
The antiferromagnetic Ising model is known to generate classical spin liquids on various flat-space lattices, such as the triangular lattice [31] and the pyrochlore lattice, the latter example being known as spin ice [32, 33] which realizes the same statistical mechanics as water ice [34]. These classical spin liquids are characterized by the absence of magnetic order and an exponentially large number of lowest-energy states, implying a finite residual entropy density.
Here we study Ising spins placed on two-dimensional hyperbolic lattices, assuming an interaction of the form (1) irrespective of the curvature of the underlying space. We demonstrate that classical spin liquids can be realized in hyperbolic space, and we elucidate on the similarities to and differences from their flat-space counterparts. A remarkable result is that the boundary shape can be used to control the low-temperature magnetic phases.
Hyperbolic tessellations and geometric frustration.—
Discretizations of curved space lead to the concept of tesselations: A uniform tessellation $\{p,q\}$ denotes a two-dimensional space-filling lattice of regular $p$-sided polygons where $q$ polygons meet at each lattice vertex. Such tessellations can naturally be embedded on two-dimensional surfaces of constant curvature $\rho$ [3, 35]. The sign of the characteristic $\chi=1/2-1/p-1/q$ distinguishes between spherical ($\chi<0$), flat ($\chi=0$) and hyperbolic ($\chi>0$) tessellations, embedded on the eponymous space. Geometric frustration requires $p$ to be odd. To generate maximal frustration, we focus on hyperbolic tessellations that are composed of triangles, $p=3$. Since the simplest realization of $\chi>0$ is given by $q=7$, we focus on the $\{3,7\}$ tessellation for most of this work.
Figure 1: Snapshot low-temperature configurations of antiferromagnetic Ising models on $\{3,7\}$ lattices. Red/blue bonds indicate parallel/antiparallel spins on the respective bond. (a) System with type-A boundary with $N=29$ spins, showing a spin-liquid configuration with antiferromagnetic alignment along the boundary. (b) Type-B boundary with $N=50$ spins, showing an ordered ferrimagnet.
Uniform hyperbolic tessellations represent infinite lattices; they can be understood as limits of finite hyperbolic uniform tilings. These finite lattices have unique features that distinguish them from their planar counterparts: The number of sites grows exponentially with linear system size, leading to strong boundary contributions. Implementing periodic boundary conditions is not straightforward: Their possible unit cells are restricted by the Gauss-Bonnet theorem and involve the solution of lattice-specific diophantine equations [3, 6]; we therefore choose to work with open boundary conditions. There are a multitude of ways to construct finite lattices, themselves leading to different shapes of the lattice boundary. We choose the iterative algorithm of Refs. 12, 35, described in more detail in the supplemental material (SM) [36]. A sample lattice is shown in Fig. 1(a) as a planar graph using Poincaré projection; we refer to this type of boundary as type-A. Other boundary shapes will be discussed below.
Thermodynamics.—
We study the thermodynamic behavior of the hyperbolic Ising model using Monte-Carlo (MC) simulations using Metropolis single-spin updates combined with parallel tempering, for details see SM [36]. Simulations for large systems are plagued by low-temperature freezing, due to the lack of multi-spin updates in our algorithm [37], such that we have restricted our simulations to systems with $N\lesssim 2000$ spins.
Results for entropy and specific heat of type-A systems are shown in Fig. 2; a snapshot low-$T$ configuration is in Fig. 1(a). The specific heat $C_{V}(T)/N$ displays a broad maximum around $T/J\approx 0.6$ whose height and shape is essentially independent of system size. At low temperatures we get enlarged statistical errors because of spin freezing. The energy $E(T)/N$, shown in the SM [36], reaches a constant upon cooling below $T/J\lesssim 0.1$; this constant approaches (J)(-J) with increasing system size, in agreement with the analytical result obtained below.
The finite-temperature corrections to the $T=0$ values of energy and specific heat appear to be of thermally activated form. This is plausible given that the smallest finite energy cost of a spin flip from a ground-state configuration in a frustrated Ising model is $2J$; this applies in particular to a finite fraction of the boundary spins, Fig. 1(a). We can fit the thermodynamic quantities accordingly, e.g.
$\frac{E(T)-E(0)}{N}=\frac{\Delta}{\gamma\exp(\Delta/T)+1}$ (2)
with the gap $\Delta=2J$ and $\gamma$ the (unknown) ratio of the number of ground and excited states.
Figure 2: Entropy (main) and specific heat (inset) per site as function of temperature for $\{3,7\}$ type-A systems. The entropy, obtained from integrating the specific heat, is used to extract a residual entropy, for details see text.
We use the specific heat, with a low-temperature fit as described, to compute the entropy according to
$S(T)=N\ln 2-\int_{T}^{\infty}\!dT^{\prime}\,\frac{C_{V}(T^{\prime})}{T^{\prime}}$ (3)
in units where $k_{B}=1$, using that $S(T\to\infty)=N\ln 2$. In practice, we integrate up to $T^{\prime}/J=100$ [38]. The results for $S(T)$ in Fig. 2 for type-A systems indicate a finite residual entropy density, $s_{\rm res}=S(T\to 0)/N$, of order $0.1$ whose dependence on system size is discussed below.
Figure 3: Uniform magnetic susceptibility $\chi$ as function of temperature $T$ for type-A systems; the inset shows $1/\chi$ at low $T$.
We also use MC to compute the zero-field uniform magnetic susceptibility $\chi(T)$, with results shown in Fig. 3. We observe a very broad maximum in $\chi(T)/N$ around $T/J=1.2$. Below $T/J\sim 0.3$ data for different system sizes strongly differ, each obeying a low-temperature Curie law, $\chi\sim 1/T$, with the prefactor decreasing with system size. As detailed below and in SM [36], the reason is that the ground-state manifold contains states with non-zero magnetization, but the magnetization distribution in this manifold has a non-trivial scaling with system size.
Correlation functions.—
As the lattices are built up in layers, it is useful to analyze the spin configurations layer by layer, where each layer can be interpreted as a one-dimensional (1D) chain. Numerical data for low-temperature NN spin correlations are shown in Fig. 4 for a moderately sized lattice, where the data have been averaged over all NN bonds of one layer (panel a) or over all NN bonds connecting two neighboring layers (panel b). In addition, thermal averages for each individual bond of a small lattice are shown in Fig. 1(b). For these type-A systems, the boundary layer is least frustrated due to the lack of outer links; at low temperatures it essentially orders into an antiferromagnetic chain pattern: depending on the system size, the boundary layer has an even (odd) number of sites, leading to zero (one) frustrated intralayer bond in the ground state, respectively; this can also be nicely seen in the snapshot shown in Fig. 1(a). Fig. 4 shows that frustration is larger for the inner layers, and correlations within the layers tend to be stronger than between layers, except for the inner core of the system. The temperature evolution of the correlations and their dependence on distance are analyzed in SM [36].
Figure 4: (a) Average NN spin correlations in each layer $L_{i}$ as function of $1/T$ of the $N=617$ type-A lattice (four layers); the fourth layer being the boundary. (b) Average NN spin correlations between neighboring layers.
Ground-state manifold.—
The manifold of lowest-energy states of the Ising model under consideration can be obtained semi-analytically, using the Fisher-Kasteleyn-Temperley (FKT) algorithm [39, 40]. For Ising models on planar graphs with odd plaquette loop lengths, there is at least one frustrated bond per plaquette. A ground-state configuration can be mapped onto a (perfect) dimer covering on the dual graph, provided that the graph can host a dimer covering at all and its number of plaquettes is even. The FKT algorithm enables to efficiently count such coverings via the skew-symmetric adjacency matrix $A^{\prime}$ of the directed dual graph, for details see SM [36]. The residual entropy is then given by
$S_{\rm res}=\ln(2)+\frac{1}{2}\ln(\det A^{\prime}),$ (4)
which we have used to compute $S_{\rm res}$ for system sizes up to $N=6000$. The corresponding numerical results for $\{3,7\}$ type-A systems are shown in Fig. 5; here we have also included data for systems with partially completed layers. MC data for $s_{\rm res}$ are shown for comparison; these are consistent with the FKT data within error bars for cases where the same system size is covered. We note that the $N=232,617$ systems feature an odd number of plaquettes such that the FKT algorithm is not directly applicable; in that case the frustrated bond on the boundary provides an additional entropy of order $\mathcal{O}(\ln(N_{\rm bnd}))$, with $N_{\rm bnd}$ the number of boundary sites. Despite significant finite-size effects, the residual entropy clearly converges to a finite value for $N\to\infty$, which we estimate as $s_{\rm res}\equiv S_{\rm res}/N=0.102(2)$.
Compared to the planar triangular-lattice Ising antiferromagnet, where single spin flips within the ground-state manifold are possible, the minimum flippable motif in the hyperbolic $\{3,7\}$ case is pair flips. This is related to the coordination number being odd.
Figure 5: Residual entropy per spin of $\{3,7\}$ type-A systems as function of inverse system size, obtained from the FKT algorithm; MC results are given for comparison.
The dimer mapping can also be used to derive the following expression for the ground-state energy [36]
$\frac{E_{GS}}{J\,N}=\frac{2\lceil N_{P}/2\rceil-N_{E}}{N},$ (5)
where the number of vertices ($N$), edges ($N_{E}$) and plaquettes ($N_{P}$) of the lattice are related by Euler’s characteristic for planar graphs, $N-N_{E}+N_{P}=1$; this differs by unity from the literature since our definition of plaquettes excludes the outside. For our hyperbolic lattice we can obtain closed expressions for $N$, $N_{E}$, and $N_{P}$ based on its iterative construction. This eventually leads to [36]
$\frac{E_{GS}}{J\,N}\approx-1+\frac{2\sqrt{5}(1+\delta_{n})}{7(1+\sqrt{5})}\left(\frac{3+\sqrt{5}}{2}\right)^{-n}$ (6)
valid for large $n$ where $n$ denotes the number of lattice layers, and $\delta_{n}$ is one if $n$ is divisible by three and zero otherwise. The ground-state energy hence approaches (JN)(-JN) for $N\to\infty$, in agreement with the numerical results in the SM [36].
Classical spin liquid.—
Our analytical and numerical data for the antiferromagnetic Ising model on type-A hyperbolic lattices are consistent with the presence of a classical spin liquid: This is indicated by the absence of a thermal phase transition, by the absence of global long-range order at $T=0$, and by a finite residual entropy density. The realization of a spin liquid in the present setting is non-trivial, as the open boundary is less frustrated than the bulk. Given that the fraction of boundary sites is finite in the $N\to\infty$ limit, $N_{\rm bnd}/N\to 0.62$ for $\{3,7\}$ type-A hyperbolic tilings [36], the system under consideration is only partially frustrated. This is reflected in the residual entropy, $s_{\rm res}=0.102(1)$, being smaller than that for the planar triangular-lattice Ising model, $s_{\rm res}=0.32306$ [31].
For comparison, we have repeated the MC simulations for the ferromagnetic Ising model on the same lattices. Here we observe clear signatures of a continuous thermal phase transition into a ferromagnetic state, consistent with earlier results [18, 19, 20, 6, 21], for details see SM [36].
Figure 6: Residual entropy per spin of $\{3,q\}$ type-A systems with $q\geq 8$ as function of inverse system size, obtained from the FKT algorithm; lines are guide to the eye.
To verify the universality of the spin-liquid behavior in the antiferromagnetic case, we have extended our study to frustrated tilings $\{3,q\}$ with $q>7$, again with type-A boundaries. Results for the residual entropy obtained from the FKT algorithm are displayed in Fig. 6, unambiguously showing a finite $s_{\rm res}$ in the large-system limit. The value of $s_{\rm res}$ decreases with increasing $q$, consistent with the fact that the fraction of (unfrustrated) boundary sites increases with $q$.
Boundary control.—
We have also studied hyperbolic $\{3,7\}$ tilings with different boundary geometry. A type-B lattice is obtained by adding one site, i.e. one outer triangle, to each boundary bond of a type-A lattice. This construction generates a family of type-B tilings, Fig. 1(b). Our numerical data for type-B systems indicate that they settle into an ordered ferrimagnetic ground state, with ferromagnetic order in each layer and antiferromagnetic alignment from layer to layer, Fig. 7. The reason is energetic: Type-B systems have more boundary bonds than type-A systems, with a zigzag boundary shape, such that perfect antiferromagnetic order along the boundary coincides with ferromagnetic intralayer order, thus tipping the balance towards ordered states. More details on the behavior of type-B systems are in the SM [36].
Figure 7: NN bond correlations $\langle\sigma_{i}\sigma_{j}\rangle$ (color-coded) for $\{3,7\}$ systems at T=0.2T\!=\!0.2. (a) Type-A with $N=85$ in the spin-liquid regime. (b) Type-B with $N=50$, indicating ferrimagnetic order.
Summary and outlook.—
We have shown that the antiferromagnetic Ising model on $\{3,q\}$ tilings realizes hyperbolic classical spin liquids, which we have characterized using thermodynamic and ground-state analyses. Our results indicate the presence of a well-defined large-system limit with open boundary conditions where the residual entropy density approaches a unique non-zero value. We have also demonstrated that spin liquidity can be controlled via the boundary: By changing the boundary shape from type-A to type-B, we can tune the energetic balance towards ordered ferrimagnetic states; the latter keep the spin-liquid manifold as low-lying excited states. Studying further boundary shapes [41] is subject of ongoing work.
Among the directions for future research are the investigation of quantum effects towards triangular hyperbolic quantum spin liquids – this suggests for instance to consider quantum dimer models on the present lattices – as well as the study of classical spin liquids in hyperbolic vector spin models. More broadly, open questions concern how lattice frustration manifests in different curved-space geometries, e.g., with black holes included [13, 42], and whether the spin liquids found here have implications for gravitational theories. We note that the finite residual entropy found here appears as an essential ingredient of the AdS/CFT correspondence: It was shown that the residual entropy of the Sachdev-Ye model [43] matches the entropy associated to the AdS2 throat arising in Reissner-Nordström charged black hole geometries [44] – a result which triggered intense research establishing holographic dualities for the Sachdev-Ye-Kitaev (SYK) model [45]. Since the Sachdev-Ye model is a spin model, our results raise the possibility that a similar duality may be realized between spin chains and antiferromagnetic Ising models. A first check here will be to compare correlation functions for spin chains at the boundary of hyperbolic spaces, as calculated e.g. in Ref. 46, with the correlation functions found here. Moreover, our results emphasize the strong influence of the boundary conditions on the ground-state degeneracy and thus on the possible presence of a gravity dual.
Acknowledgements.
We thank P. Basteiro, P. M. Cônsoli, J. Karl, R. Meyer, B. Placke, and Y. Thurn for helpful discussions. Financial support from the Deutsche Forschungsgemeinschaft through SFB 1143 (project-id 247310070), SFB 1170 (project-id 258499086), and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-id 390858490) is gratefully acknowledged.